OpenStudy (anonymous):

Solve, correct to 3 d.p ------> \[F(x)= {3 \over x}-1 ~~~~~~ x_0=1\]

OpenStudy (anonymous):

solve \[F(x)=0\]?

OpenStudy (anonymous):

Yes, for \[x_0=1\]

OpenStudy (anonymous):

why the heck don't you get x = 3 by inspection?

OpenStudy (anonymous):

Um... Not sure if it's F(x)=0... Then 3 would be easy to find.

OpenStudy (anonymous):

Oh, it's F^-1(x)

OpenStudy (anonymous):

ok i would check because if you try to solve as \[x=F(x)\] you get \[\frac{3}{x}-1=0\] \[\frac{3}{x}=1\] \[3=x\] a constant funciton and so every value of x gives 3

OpenStudy (anonymous):

\[F^{-1}(x) \]

OpenStudy (anonymous):

You find this, and then use the x_0=1

OpenStudy (anonymous):

then i guess we need \[F^{-1}(x)=\frac{3}{x+1}\] first right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then forget about it being equal to zero because the numerator is 3

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

\[\frac{3}{x}-1=y\] \[\frac{3}{x}=y+1\] \[\frac{x}{3}=\frac{1}{y+1}\] \[x=\frac{3}{y+1}\] \[F^{-1}(x)=\frac{3}{x+1}\]

OpenStudy (anonymous):

Ok, true.

OpenStudy (anonymous):

i am afraid i am not being much help here because i don't really undersand the question, but there is no way that this is equal to zero

OpenStudy (anonymous):

The answer is 1.303 if that helps in anyway?

OpenStudy (anonymous):

You have to find the root.

OpenStudy (anonymous):

then i really dont understand the question

OpenStudy (anonymous):

OK. That's fine. Thanks for helping :D